# Enhancing End Stage Renal Disease Outcome Prediction: A Multi-Sourced Data-Driven Approach

### v3

@@ 4.3 Additional Analyses
The comparison of these modeling results helps us identify the optimal observation length and the most reliable and accurate model for predicting CKD progression to ESRD. However, recognizing the limitations of focusing solely on fixed timelines, we extended our analysis to address clinical realities. In practice, the progression timeline from CKD to ESRD is often uncertain when a patient presents for evaluation. To accommodate this variability, we leveraged the selected optimal model and observation window to integrate all time-length cohorts, ranging from 6 to 30 months. This integration enables predictions of ESRD likelihood and optimal observation windows without requiring precise progression timelines, enhancing the model’s relevance and applicability in real-world clinical settings. Building on this framework, we conducted additional analyses to further refine the model, including a detailed investigation of misclassifications and their potential causes, as well as an evaluation of the impact of updated eGFR equations on racial bias in CKD predictions. These efforts aim to ensure fairness and accuracy in the model’s clinical applications.
@@ 4.3.1 Analysis of Model Misclassifications
To evaluate the performance and robustness of our optimal model, we conducted an in-depth analysis of misclassifications by examining predictive probability distributions across the entire sample set. This analysis focused on type 1 (false positive) and type 2 (false negative) errors to understand discrepancies between predicted and actual outcomes. We analyzed instances where the model incorrectly predicted ESRD progression (type 1 error) and where it failed to predict ESRD progression in patients who eventually developed the condition (type 2 error). By comparing these error types, we sought to identify common factors or patterns contributing to misclassification, thus highlighting the model’s limitations and suggesting areas for refinement.
This detailed error analysis is crucial for improving the model’s accuracy and ensuring its reliability in clinical settings. By understanding the reasons behind incorrect predictions, we can enhance the model’s ability to support clinicians in making more informed and accurate decisions regarding patient care.
@@ 4.3.2 Impact of Updated eGFR Equations on Racial Bias in CKD Predictions
The evolution of eGFR equations has aimed to address inherent limitations, particularly concerning racial bias. In 2020, the use of race-based components in eGFR calculations was scrutinized due to a lack of evidence of its biological basis and the risk of perpetuating racial stereotypes and disparities in healthcare. In response, a 2021 workgroup led by the National Kidney Foundation and American Society of Nephrology recommended an updated CKD Epidemiology Collaboration (CKD-EPI) creatinine equation 36, which excluded the race coefficient to improve accuracy and equity 37. This adjustment aimed to enhance prediction precision, especially for minority populations 38; 39.
Our analysis, using data from 2009 to 2018, applied both eGFR equations to the data. To assess the impact of the updated equation, we compared predictions between the old and new equations, focusing on improvements in accuracy and equity, particularly for minority groups.
It is important to note that adopting the updated equation may alter the cohort used for analysis, as the new eGFR values directly influence CKD stage classification. This could impact the cohort identification process, and lead to the reclassification of patients, particularly those near stage boundaries. Such changes may affect both the composition of the cohort and the model’s predictions.
@@ 5.1 Cohort for Analysis & Data Characteristics
(a) Variation in CKD stage distribution as a function of different observation window lengths.
@@
In this subsection, we present the trends observed during our data integration process and provide a detailed overview of the data characteristics for the final cohort utilized in our model performance analysis. Fig.3(a) illustrates the changes in patient distribution across CKD stages over varying observation window lengths. As the observation window extends from 6 to 30 months, there is a gradual decline in the number of patients available for analysis. This decline reflects the exclusion of patients who either lacked sufficient data within the specified window or progressed to ESRD during the period. Additionally, some patient losses may be due to mortality, but we lack the data to confirm this.
To ensure model consistency and fair comparability, we trained and tested on the same cohort. Recognizing the challenge of shrinking cohort size with extended observation windows, we focused on patients with at least 30 months of data following CKD stage 3 diagnosis and split the test data from this group. For the training data, we used the same cohort but with varying data lengths corresponding to the selected observation windows. This approach allowed us to explore the impact of different observation windows and identify the optimal one based on the performance on test data.
After establishing the cohort with at least 30 months of post-CKD stage 3 data, we compared the datasets derived from claims data only (5,518 patients), clinical data only (6,842 patients), and the merged dataset used for analysis. As shown in Fig. 3(b), merging the datasets resulted in a substantial reduction in the number of patients—to 1,422 patients with a 24-month observation window (which proved optimal in subsequent analyses) and 1,355 patients with a 30-month observation window. This reduction is mainly due to excluding patients who only had either claims or clinical data, resulting in a smaller but more comprehensive dataset for analysis.
Table 1: Data characteristics of the patient cohort for analysis
@@
Table 1 summarizes the demographic, comorbidity, claims-driven, and clinical-driven characteristics of the patient cohort, comprising 1,422 patients; 86 progressed to ESRD (6%), 1,336 did not. For patients who progressed to ESRD, the mean time from CKD stage 3 to ESRD was 4.82 ± 1.82 years. The distribution plot of progression times is provided in Supplementary Figure 1. Patients who progressed to ESRD were significantly younger ( 69.13 vs. 72.04 years, $p<0.001$ ), with notable racial disparities. African American patients were more likely to progress to ESRD (14.0% vs. 4.5%), while White patients had lower progression rates ( $81.4\%$ vs. 93.0%, $p<0.001$ ).
Comorbidities such as hypertension were prevalent in both groups (99%), but diabetes was more frequent in the ESRD group ( $73.3\%$ vs. $0.00959.0\%,p=$ 0.009 ). Secondary hyperparathyroidism and conduction/dysrhythmias were also significantly higher in ESRD patients ( $p<0.001$ and $p=0.005$ , respectively). In claims-driven features, ESRD patients showed slightly higher counts of outpatient and professional claims ( $p=0.039$ ) but no significant differences in costs. Clinical-driven features highlighted lower eGFR (17.21 vs. 22.78, $p<0.001$ ), lower hemoglobin levels (12.15 vs. $0.00114.25,p<$ 0.001 ), and higher occurrences of advanced CKD stages in the ESRD group. These differences underscore the clinical and demographic factors associated with ESRD progression.
@@ 5.2 Performance Comparison of Models Across Datasets
The results demonstrate several key findings across the different data sources. In the claims data-only scenario, deep learning models, particularly LSTM (AUROC: 0.92, F1: 0.54) and GRU (AUROC: 0.92, F1: 0.50), substantially outperformed traditional machine learning approaches such as logistic regression (AUROC: 0.72, F1: 0.33) and random forest (AUROC: 0.74, F1: 0.36). For clinical data, while the performance gap narrowed, deep learning models maintained their advantage, with LSTM achieving the highest performance (AUROC: 0.88, F1: 0.60). Most notably, the merged dataset combining both claims and clinical data yielded the best overall performance, with LSTM achieving better results across all metrics (AUROC: 0.93, F1: 0.65, AUPRC: 0.61). Comprehensive results across all models and observation windows are provided in Supplementary Table 1.
Table 2: Model performance for prediction of ESRD, using claims data-only, clinical data-only, and merged data.
@@
Model Model performance metric
@@
Logistic Regression 0.33 0.72 0.44
@@
Model Model performance metric
@@
Logistic Regression 0.54 0.76 0.47
@@
Model Model performance metric
@@
Logistic Regression 0.55 0.75 0.45
@@ 5.3 Key Feature Analysis Using Explainable AI Techniques
Having identified the 24-month observation window as optimal and LSTM as the best-performing model for ESRD prediction, we employed two complementary explainable AI techniques to understand the features driving these predictions. At the cohort level, we utilized XGBoost’s feature importance analysis to identify key predictive variables across the population. At the individual level, we applied SHAP analysis to provide detailed, patient-specific insights into model predictions.
@@ 5.3.1 Cohort-Level Key Feature Identification
Our feature importance analysis of the optimal ML model, XGBoost, revealed several key predictors for ESRD progression (see Fig.  4(a)). The most critical feature was the presence of stage 5 CKD (S5), aligning with clinical literature and expert feedback on its importance in predicting ESRD. Claims-related features, such as the number of outpatient claims (n_claims_O) and total inpatient claim expenses (net_exp_I), also ranked highly. The diverse types of top-ranked features underscore the benefit of multi-sourced data integration, enhancing the model’s predictive power. This combination of clinical and claims data not only improves the accuracy of the predictions but also supports more comprehensive and reliable decision-making in clinical practice. However, while XGBoost’s feature importance reflects the absolute contribution of each feature, it does not specify whether the influence is positive or negative.
(a) Top 30 features of importance for the XGBoost model.
(b) SHAP analysis demonstrating the variation in feature impact on ESRD risk prediction for two patients with CKD Stage 5.
Figure 4: Key Feature Analysis: Feature importance and SHAP analysis for ESRD risk prediction using a 24-month observation window.
@@ 5.3.2 Feature Impact at the Individual Patient Level Using SHAP Analysis
While the previous subsection explored feature importance at the cohort level, the uniqueness of each patient’s condition necessitates a more personalized approach. To aid healthcare providers in making tailored decisions, we applied SHAP analysis to identify the key features driving individual patient predictions.
We present SHAP force plots for two patients, Patient X and Patient Y, both of whom were correctly predicted for their risk of progressing to ESRD. Despite both having CKD Stage 5, they showed notably different patterns of contributing features, highlighting how the same clinical outcome can arise from distinct combinations of risk factors and underscoring the importance of personalized treatment strategies.
In Figure 4(b), features increasing ESRD risk are highlighted in red, while those reducing risk are in blue. For Patient X, factors such as Stage 5 CKD (S5) and a high number of outpatient claims (n_claims_DR_min) elevated the risk, while higher eGFR levels and the absence of certain markers reduced it. For Patient Y, lower risk was driven by factors such as younger age, fewer minimum inpatient claims (net_exp_O_min), and lower minimum eGFR values, despite also having a Stage 5 CKD diagnosis (S5).
This variation in key features, even among patients with the same CKD stage, highlights the importance of aiding the identification of individual differences in treatment planning. SHAP analysis offers a clear visualization of these factors, enabling healthcare providers to tailor interventions to each patient’s unique risk profile and modifiable factors.
@@ 5.4 Analysis of Model Misclassifications: Type I and Type II Errors
To explore these misclassifications, we analyzed Type I and Type II errors, as depicted in Figure 5. Notably, most incorrect predictions for patients who progressed to ESRD, yet were predicted otherwise, cluster near the lower end of the plot rather than around the 0.5 decision threshold or randomly scattered. This clustering suggests consistent factors may be influencing these errors, warranting a deeper investigation.
To understand the underlying causes of model misclassification, we conducted a detailed analysis of Type I and Type II errors. We focused particularly on Type II errors, where patients were incorrectly predicted not to develop ESRD but ultimately did progress to the condition. By comparing the 16 false negatives (patients shown in the lower part of the right plot) with the 70 true positives (patients correctly predicted to progress to ESRD, shown in the upper part of the rightplot), we aimed to identify patterns and characteristics that might explain why the model failed to identify these high-risk patients. These errors can delay essential treatment, potentially accelerating disease progression.
Conversely, Type I errors represent patients who were incorrectly predicted to develop ESRD but did not actually progress. Our analysis contrasted 17 false positives (patients shown in the upper part of the left plot) with 1,319 true negatives (patients correctly predicted not to progress, shown in the lower part of the left plot). These Type I errors are particularly concerning in clinical practice as they could lead to unnecessary medical interventions and inefficient allocation of healthcare resources, potentially diverting attention from patients at genuine risk of ESRD progression. Understanding the characteristics distinguishing these misclassified cases from correctly predicted ones may provide valuable insights for model refinement and improved clinical decision support.
Figure 5: Analysis of Model Misclassification: Type I and Type II Errors. The left plots illustrate the predicted probability distribution for all patient outcomes, highlighting the model’s performance across the dataset. The right plots focus on the incorrectly predicted cases, distinguishing between Type I errors (false positives) and Type II errors (false negatives).
Table 3: Analysis of Type I and Type II Errors: Features and Their Impact on Prediction Accuracy. This table presents representative subsets of features under various timestamps. Features in bold appear in both analyses.
@@
Based on Table 3(a), CKD Stage 5 emerges as a critical feature. For patients incorrectly predicted as not progressing to ESRD (false negatives), the mean value for CKD Stage 5 at timestamp 6 is 0.06, which is significantly lower than the 0.36 mean for those correctly predicted to progress. This suggests that these patients did not have a Stage 5 record until the subsequent timestamp. The model likely interpreted the absence of a Stage 5 indicator up to timestamp 7 as a sign that ESRD development was unlikely. However, after consulting with nephrologists and examining the literature on CKD progression, we note that some CKD patients may experience a sudden decline in kidney function, ”crashing” into dialysis, where they appear stable for a period before rapidly progressing to ESRD due to acute kidney injury or treatments during hospitalization.
Based on Table 3(b), CKD Stage 5 appears in both timestamps 6 and 7 for the incorrectly predicted group, indicating that these patients had Stage 5 records at these timestamps but did not actually progress to ESRD. This is unusual. One possible explanation could be the issue of data censoring. Our dataset spans from 2009 to 2018, so patients incorrectly predicted to develop ESRD, with continuous Stage 5 records, might have progressed to ESRD in 2019 or later. Additional data collection beyond 2018 would be needed to update and refine this analysis.
@@ 5.5 Impact of Updated eGFR Equation on Racial Bias in Predictions
(a) Incorrect predictive probabilities using the 2009 eGFR equation.
(b) Incorrect predictive probabilities using the 2021 eGFR equation.
(c) Demographic distribution patterns under the 2021 eGFR equation. Encoding mappings are: Race: 0=White, 1=Black, 2=Others; Gender: 0=Male, 1=Female; ESRD Status: 0=non-ESRD, 1=ESRD.
(d) Detailed racial distribution of patients incorrectly predicted not to develop ESRD, comparing 2009 and 2021 eGFR equations.
Figure 6: Impact of Updated eGFR Equations (2009 vs. 2021) on Prediction Errors and Racial Bias. Subplots (a) and (b) compare incorrect predictions using the two eGFR equations, while (c) and (d) explore the demographic distribution and racial disparities among incorrectly predicted patients.
When comparing predictions between the 2009 and 2021 eGFR equations (Fig  6(a),  6(b)), we observed a decrease in total incorrect predictions from 33 to 28, reducing the overall error rate from 0.0232 to 0.0228. Among these misclassifications, only 7 cases were common to both equations. Analysis of racial distribution in prediction errors (Fig  6(c),  6(d)) revealed varying patterns across different groups: incorrect predictions for white patients increased from 2 to 4, African American patients decreased from 4 to 1, and patients in the ”Other” category increased from 10 to 12. The demographic breakdown for the 2021 equation cohort, including race, gender, and ESRD outcomes, is shown in Fig  6(c). However, these numbers are clearly too small for robust insights.

### v4

@@ 4.4 Additional Analyses
Beyond the core dual-phase predictive framework outlined above, we conducted several additional analyses to enhance the robustness and fairness of our ESRD prediction models. Specifically, we investigated cases of model misclassification to identify underlying patterns and contributing factors, providing deeper insights into prediction limitations. Furthermore, we assessed the influence of recently updated eGFR equation on racial disparities within CKD progression predictions, aiming to ensure clinical fairness and improve the accuracy of our predictive models.
@@ 4.4.1 Analysis of Model Misclassifications
We examined predictive probability distributions across our sample set to analyze misclassifications from our optimal model. We evaluated both false positives (type I errors) and false negatives (type II errors) in ESRD progression predictions to identify patterns contributing to incorrect predictions. This error analysis provides insights into model limitations and highlights potential areas for refinement, ultimately improving clinical reliability and decision support.
@@ 4.4.2 Impact of Updated eGFR Equations on Racial Bias in CKD Predictions
The 2021 workgroup led by the National Kidney Foundation and American Society of Nephrology recommended an updated CKD-EPI equation that removed race coefficients, addressing concerns about racial bias in eGFR calculations 31; 32, such as race-based components lacking biological basis and risking the perpetuation of healthcare disparities 33; 34.
For our study using data from 2009-2018, we primarily used the 2009 CKD-EPI equation since this was the standard clinical calculation during the study period and informed actual physician decisions and CKD staging. We additionally applied the 2021 race-free equation in our further analysis section to compare outcomes and assess improvements in prediction accuracy and equity, particularly for minority populations. Note that adopting the updated equation may affect CKD stage classification, potentially altering cohort composition and model predictions.
@@ 5.1 Cohort for Analysis & Data Characteristics
(a) Variation in CKD stage distribution of merged data as a function of different observation window lengths.
@@
In this subsection, we present data integration trends and cohort characteristics. Fig.3(a) shows patient distribution across CKD stages with varying observation windows. Patient numbers decline as windows extend from 6 to 30 months due to insufficient data or ESRD progression within the window. The ESRD columns represent patients who progressed to ESRD after their respective observation windows. While some patient losses may be attributed to mortality, we lack confirmatory data.
To ensure model consistency and fair comparability, we trained and tested our models using the same patient cohort. Recognizing the issue of cohort sizes decreasing as the observation window lengthens, we selected patients who had at least 30 months of follow-up data after their CKD stage 3 diagnosis. We specifically chose 30 months because it represents the smallest cohort across the range of observation windows; thus, patients in this 30-month cohort are also guaranteed to be present in all shorter observation windows. For the training datasets, we utilized this same group of patients but varied the length of the data included according to different observation windows. This approach allowed us to systematically assess how different observation windows affected model performance and to identify the optimal window based on test dataset results.
After requiring $\geq$ 30 months of post–stage 3 data, we compared claim-only (n = 5,518) and clinical-only (n = 6,842) cohorts to the merged sample. Combining data reduced the cohort to 1,422 patients for the 24-month window (optimal) and 1,355 for the 30-month window (Fig. 3(b)), reflecting exclusion of those lacking one data source and yielding a more comprehensive dataset.
Table 1: Data characteristics of the patient cohort for analysis (Merged data, 24-month observation window, n=1,422). Continuous variables are shown as mean ± standard deviation and compared by independent t-test (log-transformed if non-normal); categorical variables are shown as count (percentage) and compared by chi-squared test. “Missing data (%)” indicates the proportion of missing observations for each variable. P-values are two-sided, with $p<0.05$ denoting statistical significance.
@@
Table 1 summarizes characteristics of our 1,422-patient cohort; 86 (6%) progressed to ESRD while 1,336 did not. Mean progression time from CKD stage 3 to ESRD was $4.82\pm 1.82$ years. ESRD patients were younger ($69.13$ vs. $72.04$ years, $p<0.001$) with significant racial disparities: African Americans showed higher progression rates (14.0% vs. 4.5%) while White patients had lower rates (81.4% vs. 93.0%, $p<0.001$).
Hypertension was prevalent in both groups (99%), but diabetes was more common in ESRD patients (73.3% vs. 59.0%, $p=0.009$). Secondary hyperparathyroidism and conduction disorders were also significantly higher in the ESRD group. Claims data showed slightly higher outpatient/professional claims counts for ESRD patients ($p=0.039$) but no significant cost differences. Clinically, ESRD patients had lower eGFR ($17.21$ vs. $22.78$, $p<0.001$), lower hemoglobin ($12.15$ vs. $14.25$, $p<0.001$), and more advanced CKD stages.
@@ 5.2 Performance Comparison of Models Across Datasets
The results demonstrate several key findings across the different data sources. In the claims data-only scenario, deep learning models, particularly LSTM (AUROC: 0.92, F1: 0.54) and GRU (AUROC: 0.92, F1: 0.50), substantially outperformed traditional machine learning approaches such as logistic regression (AUROC: 0.72, F1: 0.33) and random forest (AUROC: 0.74, F1: 0.36). For clinical data, while the performance gap narrowed, deep learning models maintained their advantage, with LSTM achieving the highest performance (AUROC: 0.88, F1: 0.60). Most notably, the merged dataset combining both claims and clinical data yielded the best overall performance, with LSTM achieving better results across all metrics (AUROC: 0.93, F1: 0.65, AUPRC: 0.61). Comprehensive results across observation windows are provided in Appendix E.
Table 2: Model performance for prediction of ESRD, using claims data-only, clinical data-only, and merged data (24-month observation window, n=1,422). Bold indicates the best performance within each method category (Machine Learning or Deep Learning).
@@
Model Type Model Model performance metric
@@
Machine Learning Logistic Regression 0.33 0.72 0.44
@@
Model Type Model Model performance metric
@@
Machine Learning Logistic Regression 0.54 0.76 0.47
@@
Model Type Model Model performance metric
@@
Machine Learning Logistic Regression 0.55 0.75 0.45
@@ 5.3 Key Feature Analysis Using Explainable AI Techniques
Leveraging the optimal 24-month window and our best ESRD prediction model, we performed (1) cohort-level feature importance to pinpoint the most influential predictors across the population, and (2) individual-level SHAP analysis to generate patient-specific explanations of each prediction.
@@ 5.3.1 Cohort-Level Key Feature Identification
Our feature importance analysis of the optimal ML model, XGBoost, revealed several key predictors for ESRD progression (see Fig. 4(a)). The most critical feature was the presence of CKD Stage 5 (S5), aligning with clinical literature and expert feedback on its importance in predicting ESRD. Claims-related features, such as the number of outpatient claims (n_claims_O) and total inpatient claim expenses (net_exp_I), also ranked highly. The diverse types of top-ranked features underscore the benefit of multi-sourced data integration, enhancing the model’s predictive power. This combination of clinical and claims data not only improves the accuracy of the predictions but also supports more comprehensive and reliable decision-making in clinical practice. However, while XGBoost’s feature importance reflects the absolute contribution of each feature, it does not specify whether the influence is positive or negative.
(a) Top 30 features importance for the XGBoost model. Features are colored from yellow-green (highest importance) to dark blue (lower importance).
(b) SHAP analysis demonstrating the variation in feature impact on ESRD risk prediction for two patients with CKD Stage 5. Red bars indicate features that increase predicted risk, while blue bars indicate features that decrease predicted risk.
Figure 4: Feature importance and SHAP analysis for ESRD risk prediction using XGBoost model (24-month observation window, n=1,422). See Appendix G for complete feature names.
@@ 5.3.2 Feature Impact at the Individual Patient Level Using SHAP Analysis
To support personalized decision-making, we applied SHAP analysis to quantify the features driving each patient’s risk prediction. Figure 4(b) presents SHAP force plots for two correctly predicted high-risk patients, Patient X and Patient Y, both diagnosed with CKD Stage 5 but exhibiting distinct risk profiles.
For Patient X, elevated risk is driven primarily by Stage 5 CKD (S5) and a high volume of outpatient claims (n_claims_DR_min), while higher eGFR levels and the absence of certain clinical markers mitigate that risk. In contrast, Patient Y’s lower risk contribution—despite the same Stage 5 diagnosis—stems from younger age, fewer minimum inpatient claims (net_exp_O_min), and lower minimum eGFR values.
@@ 5.4 Analysis of Model Misclassifications: Type I and Type II Errors
All subsequent analyses utilize our optimal 24-month observation window and the LSTM model, which demonstrated superior performance among all tested architectures. To explore these misclassifications, we analyzed Type I and Type II errors, as depicted in Figure 5. Notably, most incorrect predictions for patients who progressed to ESRD, yet were predicted otherwise, cluster near the lower end of the plot rather than around the 0.5 decision threshold or randomly scattered. This clustering suggests that consistent factors may be influencing these errors.
Figure 5: Analysis of Model Misclassification: Type I and Type II Errors. The left panel shows predicted probabilities for all patient outcomes (24-month conversation window, n=1,422), stratified by ESRD (green) and non-ESRD (blue) cases, with patient indices (1-1422) on the x-axis. The right panel displays only the incorrectly predicted cases (n=33), using actual patient IDs on the x-axis for identification purposes. The horizontal dashed line at 0.5 represents the classification threshold.
To understand model misclassification causes, we analyzed Type I and Type II errors. For Type II errors, we compared 16 false negatives (patients incorrectly predicted not to develop ESRD but who did) with 70 true positives (patients correctly predicted to progress). For Type I errors, we contrasted 17 false positives (patients incorrectly predicted to develop ESRD) with 1,319 true negatives (patients correctly predicted not to progress).
Table 3: Analysis of Type I and Type II Errors (24-month observation window, n=1,422): Features and Their Impact on Prediction Accuracy. This table presents representative subsets of features under various timestamps. Features in bold appear in both analyses.
@@
Based on Tables 3(a) and 3(b), CKD Stage 5 emerges as a critical feature in prediction errors. For false negatives (incorrectly predicted as not progressing to ESRD), the mean value for CKD Stage 5 at timestamp 6 is significantly lower (0.06 versus 0.36) compared to correctly predicted cases, suggesting these patients didn’t have a Stage 5 record until later. Conversely, false positives show CKD Stage 5 presence at both timestamps 6 and 7, indicating these patients had Stage 5 records but didn’t actually progress to ESRD.
@@ 5.5 Impact of Updated eGFR Equation on Racial Bias in Predictions
(a) Incorrect predictive probabilities obtained using the 2009 eGFR equation.
(b) Incorrect predictive probabilities obtained using the 2021 eGFR equation.
Figure 6: Impact of Updated eGFR Equations (2009 vs. 2021) on Prediction Errors. Subplots compare incorrect predictions using the two eGFR equations, highlighting the differences in misclassification patterns between the traditional and updated formulas.
When comparing predictions between the 2009 and 2021 eGFR equations (Fig. 6(a),  6(b)), we observed a decrease in total incorrect predictions from 33 to 28, reducing the overall error rate from 0.0232 to 0.0228. Among these misclassifications, only 7 cases were common to both equations. Analysis of racial distribution in prediction errors revealed varying patterns across different groups: incorrect predictions for white patients increased from 2 to 4, African American patients decreased from 4 to 1, and patients in the ”Other” category increased from 10 to 12. However, these numbers are clearly too small for robust insights.